if the k.e of a particle is zero ,what is its linear momentum? If the total energy of the particle is 0 ,is its linear momentum necessarily 0?

Grade:12

1 Answers

vikas askiitian expert
13 years ago

dear syed

if kinetic energy is 0 then there will be no linear velocity hence no linear momentam ...

now total energy = sum of energy possed by particle

                        =KE + PE + heat energy so on

consider if there is only potential energy & kinetic energy then

 Total energy = 0 = KE+PE

 KE = -PE   ............1

from eq 1 it can be said that if PE is equal to KE but with -ve PE then particle may have KE & hence linear momentam...

eg =>  if we throw a mass M from surface of earth to infinity then its total energy should be zero ....

TE =  PE + KE = 0

 KE = GMMe/R ............2

hence particle will have linear momentam.....